OpenStudy (anonymous):

What is the approximate length of a semicircle with a radius of 12 centimeters? A. 18.8 centimeters B. 37.7 centimeters C. 75.4 centimeters D. 150.7 centimeters

OpenStudy (***[isuru]***):

\[C = 2 \pi r\] for a semi circle it will be \[C = \pi r\](since it;s a half of a full circle ) r is the radius so... \[C = 3.14 \times 12\]\[C = 37.68 \ cm \] which is 37.7 cm hope this will help ya:)

OpenStudy (anonymous):

thanks